ms:

A cube has side length x. One side of the cube is increased by 4 inches, and another side is doubled. The volume of the new rectangular prism is 450 cubic inches. The equation 2 x cubed + 8 x squared = 450 can be used to find x. What was the side length of the original cube? Use a graphing calculator and a system of equations to find the answer.

lilianamendez2:

The equation 2 x cubed + 8 x squared = 450 = 2x^3 + 8x^2=450?

lilianamendez2:

or The equation 2 x cubed + 8 x squared = 450 => 2 x ^3 + 8 x ^2 but that make no sense - so i think no

lilianamendez2:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @lilianamendez2 The equation 2 x cubed + 8 x squared = 450 = 2x^3 + 8x^2=450? \(\color{#0cbb34}{\text{End of Quote}}\) so that mean we just solve this for the x value and calcule the x

lilianamendez2:

(2x^3 + 8x^2=450)